OpenStudy (mathlegend):

Trigonometry: Give the domain of the function f(x) = 2 / (x+4) in interval notation.

OpenStudy (mathlegend):

Totally lost...

OpenStudy (mathlegend):

@ranga can you help me with this trig? I want to do... x+4 >= 0 x >= -4 (-4, positive infinite)

OpenStudy (ranga):

This is not a trig question though.

OpenStudy (mathlegend):

Pre-Calc or Algebra?

OpenStudy (ranga):

Domain of a function is all allowed values of x. Here there is an expression in the denominator and we cannot allow it to go to 0. so x + 4 cannot be 0. or x cannot be -4. All other values of x are allowed. So the domain is (-infinity, -4) U (-4, infinity) where U is the Union symbol. Note that -4 is an open interval as indicated by the parenthesis ( and ) instead of [ and ]. Yeah, algebra.

OpenStudy (mathlegend):

Okay, I think this is a trig question but I don't think it really makes since... "Find all values of theta for which sin of theta = 0"

OpenStudy (mathlegend):

Wouldn't that just equal 0?

OpenStudy (ranga):

The sine function is a periodic function. It repeats the same cycle over and over again both in the positive and the negative direction. sin(theta) = 0 when theta = 0, pi, 2pi, 3pi ..... as well as -pi, -2pi, -3pi, .... So theta = n(pi) where n is any integer (...-3, -2, -1, 0, 1, 2, 3, ...)

OpenStudy (mathlegend):

oh wait.. so it is... 0 since Sin is y/r at point (1, 0) right?

OpenStudy (ranga):

That is the solution in the first quadrant. If they ask you for a solution in the domain [0, pi/2] then you can stop at theta = 0. But if they ask a general question with no domain restriction then you will have to say: n*pi where n is any negative, zero or positive integer.

OpenStudy (ranga):

sin(0) = sin(pi) = sin(2pi) = sin(3pi) = sin(-pi) = sin(-2pi) = sin(-3pi) = 0

OpenStudy (mathlegend):

But Sin can only be 0.. at (1,0) and (-1,0) right?

OpenStudy (mathlegend):

I am thinking of the quadrants

OpenStudy (ranga):

If they ask in the interval [0, 2pi] when does sin(theta) become zero you can say at theta = 0 and at theta = pi and stop there. If no domain restriction, then you can rotate theta = 0 by one full rotation and you will end up in the same place, so 0, 0 + 2pi, 0 + 4pi, all are the same angles.

OpenStudy (mathlegend):

So like added 360 degrees?

OpenStudy (mathlegend):

adding*

OpenStudy (ranga):

similarly theta = pi, pi + 2pi, pi + 4pi .... will all end up in the same place. A full 2pi rotation will always end up in the same location. exactly like adding 360 degrees or 2pi radians.

OpenStudy (mathlegend):

and you are saying I can add 360 each time but I can also subtract 360 each time as well to go in the opposite correction.

OpenStudy (ranga):

Yes.

OpenStudy (ranga):

Also, once you are out of basic trig where they use angles in degrees in a triangle you have to switch to thinking angle in radians in calculus. So rather than rotating by 360 degrees we will be rotating 2pi radians in calculus.

OpenStudy (mathlegend):

So I have to remember the Unit Circle?

OpenStudy (ranga):

Not necessary to memorize it. But if you do enough problems you will remember the sin and cos values in the first quadrant for the angles: 0, 30, 45, 60 and 90 degrees or 0, pi/6, pi/4, pi/3 and pi/2 radians and the rest of the angles can be derived just from the first quadrant values.

OpenStudy (ranga):

You can plot the sine curve online that provides a graphing calculator. For example, you can go to: https://www.desmos.com/calculator and type in sin(x) on the left and it will plot it for you. You can then go and click on the points where it crosses the x-axis and it will tell you the x where sin(x) is zero. As you will notice there are infinite number of x values at which sin(x) = 0.

OpenStudy (mathlegend):

@ranga how did you become so good at Math? Is this a natural talent for you?

OpenStudy (ranga):

IDK. Different people have different things that comes easily to them. I can do a couple of things okay but there are a million things I cannot do that others do very well. I cannot sing, draw or write poetry but it comes easily and naturally to some others.